OpenStudy (anonymous):

answers!!! q2 p1: Vs as vin

OpenStudy (anonymous):

@brash :can u post them here?

OpenStudy (anonymous):

q2 p1: Vs as vin<vt p2: v0 = vs-kr(e^((vin-vt)/vn)-1) p3: vin = ln(vs/2kr +1)vn+vt p4: vin=sqrt(vs/kr)+vt p5: -kre^((vin-vt)/vn) p6: -kr(vin-vt) vs=4.5 r = 1k k =1mA vt=1 vn=1 1: 4.5 vin=0.5 2: 2.7818 vin=2 3: 2.1786 vo=vs/2 4: 3.1213 vo=vs/2 mosfet 5: -2.7182 vin=2 6: -1 vin=2 mosfet all is correct

OpenStudy (anonymous):

@ brash P2 this formula didn't got me the right answer and i also tried it 4 ur specification but it also didn't get ur answer too what's wrong with that ????

OpenStudy (anonymous):

vin=2 vo = 4.5-10^3*10^-3*((e^((2-1)/1))-1) = 2,7818

OpenStudy (anonymous):

Thx @ brash - may i did something with wrong way

OpenStudy (anonymous):

still not getting yaar the formula is slight confusing. @brash can you tell me for Vs =4v and remaining as same as your questions then whats the answer.?

OpenStudy (anonymous):

part 2 and part 3 getting difficult.

OpenStudy (anonymous):

@brash , I got it all except part3 .. the formula you gave for part 3 is for finding Vin. But we need to find out Vout. How can we do that ? ... @jason10 , please see

OpenStudy (anonymous):

@brash , the problem for part 4 is also same, please help

OpenStudy (anonymous):

pls do you have the formulas for q1

OpenStudy (anonymous):

write your question text, in my question i need Vin in 3 part and Vin in 4 part after we replace ExpFET to MOSFET